OpenStudy (anonymous):

20z^2=40z+60 what are the solutions of the equation?

OpenStudy (anonymous):

Solutions: z = -1 z = 3

OpenStudy (anonymous):

you really know ur math, thank you

OpenStudy (anonymous):

@kianap really could use some clarification on other questions

OpenStudy (anonymous):

glad to help...! : )

OpenStudy (anonymous):

\[20z^2 - 40z - 60z = 0\] \[20(z^2 - 2z - 3) = 0\] \[z^2 -2z -3 = 0\] \[(z-3)(z+1) =0\] Then z = 3 and z = -1 should be your solutions. you have two solutions because two values if plugged in for z would give 0, also look up factoring quadratic equations to see how I got to the last step.